Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

DURC 10m Air Rifle Open

Distribution

Dr. Trevor West DUCAC Chairman
Aubrey Storey DUCAC Secretary
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Tony O’Donnell Student Record
Susan Thompson Trinity News

Summary

Venue UCD Rifle Club, Belfield
Date 24-25 November 2001
Discipline 10m Air Rifle

DURC Attendees

Steven Watterson Mark Dennehy
Ray Reilly  Liam Murray
Alan Smyth  Peter Byrne
Adrian Costigan  Dave Barron

Victoria McDowell

Scores

Steven Watterson 565
Mark Dennehy 554,548,533
Ray Reilly 535
Liam Murray 502
Alan Smyth 497
Peter Byrne 467,500
Adrian Costigan 456
Dave Barron 382
Victoria McDowell 363

Prizes Won

Steven Watterson Third place overall
Liam Murray First place class D
Peter Byrne Third place, Juniors

Report

The annual DURC registered 10m Air Rifle Open took place this year on the weekend of the 24th and 25th of November in the UCD Rifle Range. Attendance was high, with 35 entrants and some 43 entrys in total, from clubs as far away as Antrim. The shoot was a registered event, which means it is one of the 6 air rifle events during the year where the scores may qualify a shooter for selection to represent Ireland internationally.

The standard of competition was higher than it has been in many years, with all eight finalists scoring above 550 in the qualifying round. Two DURC members, Mark Dennehy and Steven Watterson (the secretary and captain of the club) made it to the finals. Mark placed in last place in the finals. Steven went into the finals tied for third place with Junior Squad shooter Richard Stapleton on 565 points (the olympic minimum qualifying score). A very tight finals ensued, with Richard and Steven exchanging places with nearly every shot. Steven managed to fend off Richard with his last two shots however, to take third place by 0.3 points on 662.9 points (out of a possible 709.0).

The Finals firing line. Left picture: Mark Dennehy.

Right Picture, Left to Right: Alan Lewis (Sydney Olympic Games), Steven Watterson (DURC Captain) , Richard Stapleton Jr. (Irish National Junior Squad), Rhona Barry (Atlanta Olympic Games), Nicola Murphy (Overall Winner)

John Keeney acting as Range Officer during the ten-shot Finals, with Ben Jones acting as Statistics Officer

The finals saw three such ties decided - Declan Kelly and Robert Clarke started on 561 and exchanged places several times, with Declan finally winning out by 0.7 points. Rhona Barry and Nicola Murphy both started on 574, tied for first place, but Nicola proved to be the stronger shooter on the day and won by 1.1 points in total with a final score of 673.5. Nicola was in fact the story of the day - a trigger malfunction had caused her to lose 9 points in one shot during the qualifying round - that she was able to put this disaster firmly behind her and go on to tie for first place was an inspiring performance.

Other DURC members did well also, with Liam Murray breaking 500 for the first time and claiming the first place in Class D, even though he couldn’t attend the award-giving ceremony. Peter Byrne also set a personal best, hitting 500 right on the nose. Alan Smyth has nearly returned to his old standard, just missing the 500 mark by 3 points. Ben Jones is also closing in on his previous standard, with 494. Adrian Costigan is making excellent strides in his training, with scores of 467 and 456 repectively. Newcomers Dave Barron, Steven Beck and Victoria McDowell also did admirably, though they have not reached the rostrum yet. And Ray Reilly has made a welcome return to the Air Rifle scene after a prolonged absence caused by a back injury. He put in a respectable 535.

Keith White, though he did not shoot, proved invaluable yet again as a Statistics Officer, responsible for scoring targets and tracking peoples scores during the competition, which he did using his own custom software which is due for general release soon.

Mr. Terry McAuley, the DURC president and Director of Sport for Trinity College, graciously agreed to come to the Open and present the prizes. Fortunately, he was able to observe the 10-shot Finals and was suitably impressed by the high level of shooting. He gave a short speech on the plans for the new Sports Center and it’s planned Rifle Range, and then presented the prizes. We’d like to thank Mr. McAuley for donating his time to the club for this event, as it certainly made an impression on the target shooting community.

We would also like to thank Hammerli, who sponsored us with some baseball caps for all the Junior shooters at the event, and especially Mr. Simon Kennedy, an ex-Captain of DURC, who gave a large contribution to the prize fund. As Simon was a member of the National Air Rifle Squad for many years and represented Ireland internationally in Air Rifle, we felt that the commissioning of the Simon Kennedy Air Rifle Shield was appropriate: it will be awarded each year to the overall winner of the DURC 10m Air Rifle Open.

Director of Sport Terry McAuley presents the Simon Kennedy Air Rifle Shield to the overall winner Nicola Murphy, pictured here with Runner-up Rhona Barry and bronze medallist Steven Watterson.

Misé le meas,

Mark Dennehy, DURC Secretary